OpenStudy (anonymous):

Find the solution to the differential equation \[\frac{ dy }{ dx }+\frac{ y }{ 5 }=0 \] with the initial condition y(0)=8

OpenStudy (anonymous):

So, ive done separation of variables, and integrated dy/(-y) amd dx/5 and gotten -ln(y) = x/5+C, but I am not sure what to do next.

hartnn (hartnn):

try to isolate y

hartnn (hartnn):

if ln a= b then a=e^b

OpenStudy (anonymous):

when x=0 y=8 plug that solve for c and yeah isolate y

OpenStudy (anonymous):

ok, so \[-\ln(y)=\frac{ x }{ 5 }+C\] becomes \[-y=e^(\frac{ x }{ 5 }+C)\]

OpenStudy (anonymous):

so -8=e^(0+c), -8=e^c?

OpenStudy (anonymous):

c=-ln(8) ?

hartnn (hartnn):

yup, c=-ln 8 so -ln y +ln 8 = x/5 ln (y/8) = -x/5 got this ?

hartnn (hartnn):

property of log, log A-log B =log (A/B)

OpenStudy (anonymous):

ahhhh yes I see that

OpenStudy (anonymous):

y=8*e^(x/5)

hartnn (hartnn):

y=8*e^(-x/5)

OpenStudy (anonymous):

Haha yes, making every mistake posssible on this one. Thanks for the help!

hartnn (hartnn):

welcome ^_^